You pay your secretary $6.75 per hour. A stamped envelope costs 42 cents, and regular stationery costs 5 cents per page, butPart c, d and e
Answer #1

a). A 9-page letter would have its 1st page as the fancy letter head and 8 pages of regular stationery.

Therefore, the stationery cost alone for a 9-page letter   is 1*0.16+8*0.05 = $ 0.56.

(b). If the secretary spands 4 hours on typing and corrections, then the amount payable to the secretary is 4* $ 6.75 = $ 27.00. On adding the cost of the envelope and the stationery, the total cost of preparing and mailing the letter is $ 0.56 + $ 0.42+$ 27 = $ 27.98.

( c). If c is the cost in dollars of the stationery and if p is the number of pages, then c(p) = 0.16+(p-1)*0.05 = 0.16+0.05p – 0.05 = 0.05p+0.11. Thus, c(p) = 0.05p+0.11.

( d). Let h be the number of hours spent by the secretary in preparing a p-page letter. Then c(h,p) = 6.75h + 0.42+ 0.05p+0.11= 6.75h+0.05p + 0.53. Thus, c(h,p) = 6.75h+0.05p + 0.53.

(e). The cost of preparing and mailing a 4-page letter which took the secretary 35 minutes of prepare, is 6.75*(35/60)+0.05*4+0.53 = 3.94 +0.20+0.53 = $ 4.67.